ES016A

PI: Spencer
Title: SS433

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A/dual pol
Obs time (UT): EU 04:00(15/02)-15:00(15/02)
Stations: EbMcNtOnTrWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Sat Feb 15 10:21:11 EST 1997) Minor failures
Scans from 5:19 to 5:38 and from 9:27 to 9:49 lost due to tape problems.

Onsala (John Conway at Sat Feb 15 10:56:41 EST 1997) Minor failures
Problem with schedule, check procedure every scan but no time allowed, gave late tape start so lose first third of the 2 min long passes, until around 9:30UT when check procedures were removed. No preob, tsys not implemented in postob,so sparse Tsys before 09:30Ut, afterwards Tsys implemented in postob every scan. Last 30 minutes sources below elevation limits (6 deg)

Onsala (John Conway at Sat Feb 15 10:56:57 EST 1997) Minor failures
Problem with schedule, check procedure every scan but no time allowed, gave late tape start so lose first third of the 2 min long passes, until around 9:30UT when check procedures were removed. No preob, tsys not implemented in postob,so sparse Tsys before 09:30Ut, afterwards Tsys implemented in postob every scan. Last 30 minutes sources below elevation limits (6 deg)

Effelsberg (dgraham at Mon Feb 17 02:49:48 EST 1997) Minor failures
Formatter probably 1 sec slow. No Tsys since no time for tsys allocated.

Westerbork (Tony Foley at Tue Feb 18 10:03:03 EST 1997) Success
Looks OK

Cambridge (Alastair Gunn at Wed Feb 19 09:09:19 EST 1997) Minor failures
No major problems. Field system crashes at 07:21 (restarted for 07:36), 08:05 (restarted for 09:21 due to repeated crashes), 11:25 (restarted for 11:32). Weather windy, cloudy. Tapes JODV0016,D0F1 JODR0139,7B40 and JODR0192,E77D sent to Bonn via Woodcroft on 17/02/97.

Torun (Roman Feiler at Wed Feb 19 16:11:08 EST 1997) Minor failures
BBC07 failed, excluded from procedures, from 13:46 sources elevation too small, schedule halted.

Jodrell (Alastair Gunn at Fri Feb 21 03:28:41 EST 1997) Minor failures
POLARISATIONS ARE CROSSED. No other known problems. Weather cloudy and windy. Tapes EVN00189,6691 EVN00131,6ACD and IRAM0137,7FD3 sent to Bonn via Woodcroft on 17/02/97.

Calibration (Fredrik T Rantakyro at Thu May 8 11:30:24 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name es016a.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Tue Sep 23 14:40:56 METDST 1997) Minor failures
Torun tapes unreadable at correlator.